OpenStudy (anonymous):

two similar pyramids have lateral areas of 20 ft^2 and 45 ft^2. the volume of the smaller pyramid is 8ft^3. what is the volume of the larger pyramid

jhonyy9 (jhonyy9):

20 --- 8 45 --- x ------- x= (8)(45)/20 = 360/20 = 18 - the volume of the larger pyramid is 18 ft*3